# Extending Nonlocal Kinetic Energy Density Functionals to Isolated Systems via a Density-Functional-Dependent Kernel

### v2

@@ Abstract
Conclusion—In summary, we systematically demonstrate that the ill-defined average charge density $\rho_{\mathrm{avg}}$ in isolated systems induces the BC instability by violating scaling laws and disrupting Pauli energy positivity. By introducing a density-functional-dependent kernel, we eliminate this instability while preserving the computational efficiency and UEG accuracy of the original WT framework. In 56 single-atom calculations, ext-WT reduces total energy errors by an order of magnitude compared to WT KEDF, while maintaining the lowest charge density errors among all tested functionals. This work establishes a generalizable strategy for designing nonlocal KEDFs that bridge bulk and isolated systems, addressing a longstanding limitation in OFDFT.
Acknowledgements.
The work of L.S. and M.C. was supported by the National Key R&\&&D Program of China under Grant No. 2025YFB3003603 and the National Natural Science Foundation of China under Grant No. 12135002. The numerical simulations were performed on the High-Performance Computing Platform of CAPT.

### v3

@@ III Results
IV Conclusions
In summary, we systematically demonstrate that the ill-defined average charge density $\rho_{\mathrm{avg}}$ in isolated systems induces the BC instability by violating scaling laws and disrupting Pauli energy positivity. By introducing a density-functional-dependent kernel, we eliminate this instability while preserving the computational efficiency and UEG accuracy of the original WT framework. In 56 single-atom calculations, ext-WT reduces total energy errors by an order of magnitude compared to WT KEDF, while maintaining the lowest charge density errors among all tested functionals. This work establishes a generalizable strategy for designing nonlocal KEDFs that bridge bulk and isolated systems, addressing a longstanding limitation in OFDFT.
V Acknowledgements
The work of L.S. and M.C. was supported by the National Natural Science Foundation of China Excellence Research Group Program No. 12588301 and the National Key R&\&D Program of China under Grant No. 2025YFB3003603. The numerical simulations were performed on the High-Performance Computing Platform of CAPT. This work was partially supported by PKU Kunpeng&Ascend Center of Excellence.